OpenStudy (anonymous):

A cup of coffee is poured at 190F, and is allowed to cool in a 70F room. The temperature of the coffee is 170F after half an hour. How long will it take thte coffee to cool to 120?

OpenStudy (anonymous):

exponential decay?

OpenStudy (anonymous):

190 = ae^kt + 70 190 = a + 170 a = 120 170 = 120e^k30 + 70 100 = 120e^k30 100/120 = e^30k ln(100/120) = 30k ln (100/120)/30 is approx -.0060773852 (do I need to have that exact of a value?)

OpenStudy (anonymous):

lebron james

OpenStudy (anonymous):

Newton's law of cooling

OpenStudy (anonymous):

no wait its kobe bryant

OpenStudy (anonymous):

So then I got T(t) = 120e^-0.0060773852 + 70 120 = 120e^-0.0060773852t+70 50/120 = 120 e^-0.0060773852t ln(50/120) / -0.0060773852 is approx 144 minutes

OpenStudy (anonymous):

kevin durant

OpenStudy (anonymous):

@batman35, can you stop replying? you are not helping

OpenStudy (anonymous):

does it look right?

OpenStudy (anonymous):

i'll look it over..

OpenStudy (anonymous):

shouldn't it be like this? p(t) = p_o * e^(-kt)+70 170 = 190*e^(-k*.5)+70

OpenStudy (anonymous):

on which part?

OpenStudy (anonymous):

to find 'k'

OpenStudy (anonymous):

yeah, I wasn't sure if I had the right numbers

OpenStudy (anonymous):

i'm not sure what newton's law of cooling is, but i am familiar with exponential decay..

OpenStudy (anonymous):

how does that change my k value?

OpenStudy (anonymous):

190 would be the initial temperature of the object

OpenStudy (anonymous):

@ParthKohli @Hero

OpenStudy (anonymous):

@campbell_st @SolomonZelman

OpenStudy (loser66):

@jtryon while DemolisionWolf is helping, please don't tag others. Please respect the helper who is trying helping you.

OpenStudy (anonymous):

I only did that because it seems like he wasn't sure

OpenStudy (anonymous):

Do I wait for the helper to tell me that he doesn't know and then I can tag others?

OpenStudy (anonymous):

hold on, so what is your quesiton? are u asking if you did it correct?

OpenStudy (anonymous):

I'm asking if I did it right or I put the numbers in the wrong places because I don't know if 144 minutes is right

OpenStudy (anonymous):

see how you started here: 190 = ae^kt + 70 a actually means beginning temperature, so its like this: T(t) = 190 e^(-k*t) + 70 -k because it is decaying, positive k would be increasing

OpenStudy (anonymous):

so that is going to change my answer

OpenStudy (anonymous):

but don't you plug in the value for a so you have 170 = 120e^k30+70?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (loser66):

the equation of the Newton's cooling Law \[T(t) = T_{room}+ (T_{initial}-T_{room})e^{-kt}\] so that your equation is \[T(t) = 70 + 120e^{-kt}\] to me, your work is ok until it Now, after 30 minutes, the temperature of the coffee is 170 It means T(30) = 170, plug into the formula to find k \[T(30) = 70 +120 e^{-30k}= 170\\100 = 120 e^{-30k}\\\dfrac{100}{120}=e^{-30k} =\dfrac{10}{12}\] \[e^{30k}=\dfrac{12}{10}\rightarrow 30k = ln\dfrac {12}{10}\rightarrow k = 0.006\] replug into the formula \[T(t) = 70+120e^{-0.006t}\] we need to find out how long will it take to cool it down to 120 that means \[120 = 70 +120e^{-0.006t}\\\frac{50}{120}=e^{-0.029t}\rightarrow ln\dfrac{12}{5}=0.006t \\t =145.9 ~min\] that's all I know.

OpenStudy (anonymous):

@Loser, where did you get -0.029t?

OpenStudy (anonymous):

I see where you got -0.006t

OpenStudy (anonymous):

should it not be 0.006 instead of -0.006? I am trying to follow along with the way that you worked out the problem

OpenStudy (anonymous):

because k -> 0.006

OpenStudy (loser66):

nope, typo, not 0.029, it is0.006

OpenStudy (anonymous):

and it is not -0.006 but 0.006?

OpenStudy (anonymous):

so you get ln (12/5) / .006 is approximately 146 mins

OpenStudy (anonymous):

don't know why you put -30k and - as you were going through it

OpenStudy (loser66):

yup

OpenStudy (anonymous):

the constants to me seem like they are +30k and +0.006

OpenStudy (loser66):

Actually, I made a short research to know what the Newton Cooling law is and worked on your problem as above. ( since I saw no one helped you hihihihi) However, I don't think I make any mistake except the typo as corrected. Hopefully it is correct. Good luck

OpenStudy (anonymous):

It is correct. I was just trying to figure out where -30k came from since 30 represents the time and it is positive

OpenStudy (loser66):